# From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?

### v2

@@ 4.1. Topic Diversity (RQ1 & RQ2)
News | Post | News Post | |
@@
0.179 | ✗ | ✗ |
Notably, as observed in Fig. 3(b) and the coefficient estimates of in Columns (1) and (2), the difference between true and false news posts is larger than that between true and false news claims. Additionally, the coefficient estimate of becomes significantly positive ( , ), which suggests that false news posts containing conspiracy theories exhibit even higher topic diversity than false news posts without conspiracy theories. There are two possible reasons for the larger difference in topic diversity between true and false news at the post level compared to the claim level. One is that news claims with higher topic diversity receive more posts on X. Another is that false news has a bigger increase in topic diversity from the claim level to the post level, compared to true news. Given the second reason, we further examine the change in topic diversity from the source news claims to the corresponding posts. In Column (3) of Table 5, the positive coefficient estimate of is statistically significant ( , ). It suggests that the increase of topic diversity in false news from the claim level to the post level is bigger than that in true news, which could partially explain the enlarged difference in topic diversity between true and false news posts, compared to their source claims. In addition, the coefficient estimate of in Column (3) of Table 5 is also significantly positive ( , ). It means that the topic diversity in false news that contains conspiracy theories increases more than that in false news without conspiracy theories from the claim level to the post level. This can explain why the coefficient estimate of in Column (1) of Table 5 is statistically insignificant, whereas it becomes statistically significant in Column (2).
4.2. News Sharer Engagement: Lifetime and Post Count (RQ3 & RQ4)
Summary statistics: Fig. 3(d) shows that false news items have significantly longer lifetimes (mean of 1463.757 hours) than true news items (mean of 47.951 hours, , ). CT news items have longer lifetimes (mean of 1968.141 hours) than false news items ( , ). Additionally, Fig. 3(e) shows that false news claims receive more posts (mean of 38.015) than true news claims (mean of 3.159, , ). However, the number of posts received by CT news claims (mean of 44.976) has no statistically significant difference from that received by false news claims ( , ).
Regression results: As shown in Column (1) of Table 6, the coefficient estimate of is significantly positive ( , ), which suggests that false news items have longer lifetimes to get new posts on X than true news items. Additionally, the coefficient estimate of is also significantly positive ( , ). It means that false news items containing conspiracy theories have significantly longer lifetimes than false news items without conspiracy theories. However, the coefficient estimates for and are statistically insignificant. This suggests that topic diversity has no significant impact on the lifetime of news items, irrespective of the veracity.
Additionally, in Column (2) of Table 6, the coefficient estimate of is significantly positive ( , ), indicating that the number of posts referencing a specific source news claim is significantly associated with its veracity. On average, false news claims are linked by times more posts on X compared to true news claims. Notably, the coefficient estimate of is not statistically significant for post count, which implies that the integration of conspiracy theories has no additional impact on the number of posts received by false news claims. The coefficient estimate of is statistically insignificant. This suggests that the topic diversity of a source news claim has no significant relationship with the number of online posts it receives. The reason for the enlarged difference in topic diversity between true and false news posts is not supported by this observation. Furthermore, the coefficient estimate for the interaction is also statistically insignificant, reinforcing the notion that topic diversity has no significant effect on post count, irrespective of the veracity of news claims.
@@ 4.2. News Sharer Engagement: Lifetime and Post Count (RQ3 & RQ4)
Lifetime | Post Count | |
@@
0.481 | 0.188 |
4.3. Post Viewer Engagement: Reposts, Likes, and Replies (RQ3 & RQ4)
Summary statistics: Figs. 3(f) – 3(h) reveal that, on average, true news posts get more engagement on X than false news posts in terms of reposts (true: 29.396, false: 14.403, , ), likes (true: 84.765, false: 45.234, , ), and replies (true: 6.294, false: 5.031, , ). However, it’s worth noting that the news posts that receive the most significant reposts, likes, or replies belong to the false category. Additionally, CT news posts receive more reposts (mean of 37.942, , ), likes (mean of 106.004, , ), and replies (mean of 25.173, , ) than overall false news posts. Notably, the difference in reply count between true and false news posts is the smallest compared to repost count and like count.
Regression results: In Column (1) of Table 7, the coefficient estimate of is significantly negative ( , ), which means that false news posts receive fewer reposts than true news posts. On average, false news posts receive 22.6% fewer reposts than true news posts. Similarly, the coefficient estimate of in Column (2) is significantly negative ( , ). On average, false news posts receive 14.4% fewer likes than true news posts. However, the coefficient estimate of in Column (3) is significantly positive ( , ). This means that, on average, false news posts receive 49.2% more replies than true news posts. Together with the observations in Figs. 3(f) – 3(h), the results indicate that the interaction of reply with the veracity of posts may vary from the interactions of reposting and liking.
Additionally, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive. It indicates that heightened topic diversity is associated with increased online engagement in terms of reposts, likes, and replies. However, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly negative. This suggests that the positive association of topic diversity with engagement on X is reduced within the context of false news posts. On the contrary, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive. This means that the integration of conspiracy theories is positively linked to the engagement with false news posts on X at post viewer level. Specifically, false news posts that contain conspiracy theories receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
@@ 4.3. Post Viewer Engagement: Reposts, Likes, and Replies (RQ3 & RQ4)
4.4. Robustness Checks
To ensure the robustness of our findings, we conduct several additional checks. (i) The cross-correlations among the independent variables are reported in Table S1 (news level) and Table S2 (post level). The variance inflation factors (VIFs) are also reported in Table S3. All the VIFs are significantly lower than the threshold of 4, which indicates that multicollinearity is not an issue for our analysis. (ii) We recalculate topic diversity at the theme level and repeat our analysis (Table S4 – Table S6). (iii) We further balance the samples between true and false news items through propensity score matching and reduce potential unobserved confounding factors (Table S7 – Table S9). (iiii) We use an alternative topic modeling method (BERTopic) to cluster topics again and repeat our analysis (Table S10 – Table S12). All the results are robust and consistently support our main findings.
5. Discussion
The spread of misinformation surrounding health crises is a serious public concern. In this paper, we present a comprehensive analysis of the impact of topic diversity and conspiracy theories on the engagement dynamics of COVID-19 misinformation. Specifically, we examine user engagement at two levels spanning from the origin as source news items to their propagation through online posts and subsequent social engagement (i. e., reposts, likes, and replies) on X. Our analysis yields several key findings: (i) False news, especially when accompanied by conspiracy theories, exhibits higher topic diversity compared to true news. This discrepancy increases from source news claims to news posts on X. (ii) At news sharer level, neither the news lifetime nor the number of news posts is significantly associated with topic diversity. However, false news has a longer lifetime and receives more posts compared to true news. Additionally, the integration of conspiracy theories is associated with a longer lifetime for false news but has no significant association with the number of false news posts. (iii) At post viewer level, posts with higher topic diversity receive more reposts, likes, and replies. However, the positive effect of topic diversity is moderated by the news veracity, with higher topic diversity being linked to more engagement with true news posts compared to false news posts. (iiii) The integration of conspiracy theories is linked to more engagement with COVID-19 misinformation at post viewer level. False news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. These findings have useful implications for both research and practical applications.
Implications: First, the observed high topic diversity in COVID-19 misinformation indicates that the interconnection of multiple topics becomes a characteristic of misinformation narratives during health crises. Previous research also reveals a coordinated effort among misinformation websites from multiple domains to hyperlink to each other (Hanley et al., 2023). Notably, the topic diversity in COVID-19 misinformation significantly increases from its source claims to online posts. This mutation may contribute to the adaptability of misinformation, allowing it to resonate with diverse audiences and contexts (Yan et al., 2022). Additionally, the finding that misinformation containing conspiracy theories exhibits the strongest topic diversity aligns with previous research (Miani et al., 2022), highlighting the interconnected nature of conspiracy narratives. Within the context of multi-topic misinformation, individuals tend to adaptively connect related information and make associative inferences to reduce uncertainty and gain control in a way that is congruent to their prior knowledge (Lee et al., 2023; Miani et al., 2022). Understanding this cognitive process is essential for developing targeted interventions aimed at reducing susceptibility to multi-topic misinformation, especially when accompanied by conspiracy theories.
Second, at the news sharer level, the observed longer lifetime and higher post count on X for false news items, compared to true news items, indicate the endurance and prevalence of misinformation on social media during COVID-19. Notably, few prior studies focus on engagement at news sharer level and study how false news enters the social media platforms via news sharers. The extended lifetime we observe in this paper may provide false news with more opportunities to accumulate engagement on social media, potentially reaching a broader audience. Moreover, the lifetime and post count appear to be largely attributed to the veracity of the news claims, as opposed to other variables including and . Given this, it is plausible that the prolonged lifetime and high post count for false news claims mainly result from deliberately planned dissemination strategies by misinformation providers, involving means such as social bots and monetization (Teng et al., 2022; Moffitt et al., 2021; Ballard et al., 2022). Consequently, the role of individuals’ self-motivated sharing behaviors in the endurance and prevalence of COVID-19 misinformation is limited at news sharer level. However, massive posts repeating the same false claim can still increase individuals’ susceptibility and beliefs. In this regard, social media platforms should take measures to restrict repeated original posts from the same false claims and their mutations, which is a lack of discussion in the previous research.
Third, as opposed to the engagement at news sharer level, the engagement at post viewer level (i. e., reposts, likes, and replies) on X might be largely driven by human users rather than social bots during COVID-19 (Teng et al., 2022). Organic user engagement from human users can be influenced by multiple factors. For instance, our results in Table 7 show that the engagement at the post viewer level can be affected by factors related to content (e. g., topic diversity, conspiracy theories, emotions) and context (e. g., verified status and follower count). Previous findings on the virality of misinformation at post viewer level are mixed (Vosoughi et al., 2018; Drolsbach and Pröllochs, 2023). We also find that false news posts receive fewer reposts and likes but receive more replies, compared to true news posts. Of note, the veracity of information is not explicit to users before being verified. Therefore, there might be other content or context factors that are related to the veracity and foster engagement. While a large body of research has examined the engagement at the posts viewer level (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b, a; Wang et al., 2023a; Solovev and Pröllochs, 2022), the understanding on the role of topic diversity and conspiracy theories in shaping engagement of post viewers with misinformation is still missing. Our findings show that higher topic diversity is associated with more reposts, likes, and reposts for both true and false COVID-19 news posts on X. However, despite COVID-19 misinformation having higher topic diversity, the positive effect of topic diversity on its engagement is reduced, requiring further detailed study. Notably, conspiracy theories can significantly enhance engagement with COVID-19 misinformation. To efficiently reduce social engagement with (health) misinformation, social media platforms should exclusively focus on the typical misinformation that exhibits high topic diversity and contains conspiracy theories.
Limitations: Our empirical study reports associations between variables and avoids making causal claims. Future research could benefit from experimental designs to establish causal relationships and better understand the influence of topic diversity and conspiracy theories on engagement with online misinformation during health crises. Moreover, our analysis relies on data from X, limiting the generalizability of our findings to other social media platforms. Though the prevalence of health misinformation was the highest on X during COVID-19 (Suarez-Lledo and Alvarez-Galvez, 2021), different platforms may exhibit unique engagement patterns and user behaviors, necessitating further research across the platforms to capture a comprehensive understanding of online misinformation dynamics during health crises. Additionally, the absence of detailed engagement timelines at post viewer level in our dataset hinders a more granular analysis of how topic diversity and conspiracy theories influence engagement with COVID-19 misinformation over time on social media.
6. Conclusion
The spread of health misinformation on social media, especially during health crises such as COVID-19, poses unprecedented threats to our society. It is crucial to understand the factors and mechanisms that contribute to the engagement with health misinformation on social media. Here, we provide a comprehensive analysis of how topic diversity and conspiracy theories shape engagement with COVID-19 misinformation from news sharer level to post viewer level. We find that high topic diversity and the integration of conspiracy theories are characteristics of COVID-19 misinformation. Topic diversity has no significant association with the engagement at news sharer level. However, topic diversity and conspiracy theories can significantly enhance engagement with COVID-19 misinformation at post viewer level. Our findings provide insights into understanding the engagement with health misinformation and highlight the engagement patterns for news sharers and post viewers regarding topic diversity and conspiracy theories. These insights are valuable for developing targeted interventions at both user levels during health crises.
7. Ethics Statement
Acknowledgements.
- (1)
- Ballard et al. (2022) Cameron Ballard, Ian Goldstein, Pulak Mehta, Genesis Smothers, Kejsi Take, Victoria Zhong, Rachel Greenstadt, Tobias Lauinger, and Damon McCoy. 2022. Conspiracy brokers: Understanding the monetization of YouTube conspiracy theories. In WWW.
- Chen et al. (2023) Lucia L. Chen, Steven R. Wilson, Sophie Lohmann, and Daniela V. Negraia. 2023. What are you anxious about? Examining subjects of anxiety during the COVID-19 pandemic. In ICWSM.
- Chuai et al. (2022) Yuwei Chuai, Yutian Chang, and Jichang Zhao. 2022. What really drives the spread of COVID-19 tweets: A revisit from perspective of content. In DSAA.
- Chuai et al. (2024) Yuwei Chuai, Haoye Tian, Nicolas Pröllochs, and Gabriele Lenzini. 2024. Did the roll-out of Community Notes reduce engagement with misinformation on X/Twitter?. In CSCW.
- Chuai et al. (2023) Yuwei Chuai, Jichang Zhao, Nicolas Pröllochs, and Gabriele Lenzini. 2023. Political Elites in False Statements on the Internet. ArXiv.
- Drolsbach and Pröllochs (2023) Chiara Drolsbach and Nicolas Pröllochs. 2023. Diffusion of community fact-checked misinformation on Twitter. In CSCW.
- Falkenberg et al. (2022) Max Falkenberg, Alessandro Galeazzi, Maddalena Torricelli, Niccolò Di Marco, Francesca Larosa, Madalina Sas, Amin Mekacher, Warren Pearce, Fabiana Zollo, Walter Quattrociocchi, and Andrea Baronchelli. 2022. Growing polarization around climate change on social media. Nature Climate Change (2022), 1–8.
- Freiling et al. (2021) Isabelle Freiling, Nicole M Krause, Dietram A Scheufele, and Dominique Brossard. 2021. Believing and sharing misinformation, fact-checks, and accurate information on social media: The role of anxiety during COVID-19. New Media & Society (2021), 14614448211011451.
- Hanley et al. (2023) Hans W. A. Hanley, Deepak Kumar, and Zakir Durumeric. 2023. A golden age: Conspiracy theories’ relationship with misinformation outlets, news media, and the wider Internet. In CSCW.
- Kim and Kim (2023) Soojong Kim and Jisu Kim. 2023. The information ecosystem of conspiracy theory: Examining the QAnon narrative on Facebook. (2023).
- Kou et al. (2017) Yubo Kou, Xinning Gui, Yunan Chen, and Kathleen Pine. 2017. Conspiracy talk on social media: Collective sensemaking during a public health crisis. In CSCW.
- Lee et al. (2023) Sian Lee, Haeseung Seo, Dongwon Lee, and Aiping Xiong. 2023. Associative inference can increase people’s susceptibility to misinformation. In ICWSM.
- Papakyriakopoulos and Goodman (2022) Orestis Papakyriakopoulos and Ellen Goodman. 2022. The impact of Twitter labels on misinformation spread and user engagement: Lessons from Trump’s election tweets. In WWW.
- Prochaska et al. (2023) Stephen Prochaska, Kayla Duskin, Zarine Kharazian, Carly Minow, Stephanie Blucker, Sylvie Venuto, Jevin D. West, and Kate Starbird. 2023. Mobilizing manufactured reality: How participatory disinformation shaped deep stories to catalyze action during the 2020 U.S. presidential election. In CSCW.
- Recuero et al. (2020) Raquel Recuero, Felipe Bonow Soares, and Anatoliy Gruzd. 2020. Hyperpartisanship, disinformation and political conversations on Twitter: The Brazilian presidential election of 2018. In CSCW.
- Schmitz et al. (2023) Matheus Schmitz, Goran Muric, and Keith Burghardt. 2023. Detecting anti-vaccine users on Twitter. In ICWSM.
- Sharma et al. (2022) Karishma Sharma, Yizhou Zhang, and Yan Liu. 2022. COVID-19 vaccine misinformation campaigns and social media narratives. In ICWSM.
- Solovev and Pröllochs (2022) Kirill Solovev and Nicolas Pröllochs. 2022. Moral emotions shape the virality of COVID-19 misinformation on social media. In WWW.
- Tang et al. (2023) Chuang Tang, Shaobo (Kevin) Li, Yi Ding, Ram D. Gopal, and Guanglei Zhang. 2023. Racial discrimination and anti-discrimination: The COVID-19 pandemic’s impact on Chinese restaurants in North America. Information Systems Research (2023).
- Teng et al. (2022) Xian Teng, Yu-Ru Lin, Wen-Ting Chung, Ang Li, and Adriana Kovashka. 2022. Characterizing user susceptibility to COVID-19 misinformation on Twitter. In ICWSM.
- Wang and Rzeszotarski (2023) Luping Wang and Jeffrey M. Rzeszotarski. 2023. Understanding motivational factors in social media news sharing decisions. In CSCW.
- Wang et al. (2023a) Yuping Wang, Chen Ling, and Gianluca Stringhini. 2023a. Understanding the use of images to spread COVID-19 misinformation on Twitter. In CSCW.
- Wang et al. (2023b) Ying Wang, Hongwei Zhang, Wang Gao, and Cai Yang. 2023b. Spillover effects from news to travel and leisure stocks during the COVID-19 pandemic: Evidence from the time and frequency domains. Tourism Economics 29, 2 (2023), 460–487.
- Yan et al. (2022) Muheng Yan, Yu-Ru Lin, and Wen-Ting Chung. 2022. Are mutated misinformation more contagious? A case study of COVID-19 misinformation on Twitter. In WebSci.
Supplementary Materials
@@ Acknowledgements.
We additionally examine the cross-correlations and variance inflation factors among the independent variables to mitigate the concern about multicollinearity. Table S1 reports the cross-correlations among the variables in news items. Table S2 reports the cross-correlations among the variables in posts. The cross-correlations in in news items and posts are fairly small. Additionally, the variance inflation factors (VIFs) are reported in Table S3. We find that the variance inflation factors for the independent variables are all close to one and well below the critical threshold of four, which indicates that multicollinearity is not an issue of our analysis.
@@
We recalculate topic diversity at the theme level and repeat our analysis. Table S4 reports the estimation results for the topic diversity. Table S5 reports the estimation results for the engagement at news sharer level. Table S6 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
News | Post | News Post | |
@@
Lifetime | Post Count | |
@@
To reduce potential unobserved confounding factors and ensure the robustness of our results, we further conduct propensity score matching for true and false news claims. Specifically, we match true and false news claims based on the variables of emotions and the number of words, with the caliper of 0.3, “common support”, and “noreplacement” settings. Consequently, we discard 30 false news claims and achieve all biases lower than 0.05. Subsequently, we repeat our analysis based on the match true and false news claims and their associated posts on X. Table S7 reports the estimation results for the topic diversity. Table S8 reports the estimation results for the engagement at news sharer level. Table S9 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
News | Post | News Post | |
@@
Lifetime | Post Count | |
@@
To ensure the generalization of our results across the topic modeling methods, we use BERTopic, a popular unsupervised topic modeling tool, as an alternative method to automatically cluster topics.333https://maartengr.github.io/BERTopic/index.html The BERTopic model generates topic clusters for the whole news claims and posts. We use the topic clusters and their corresponding probabilities to calculate topic diversity again and repeat the analysis. Table S10 reports the estimation results for the topic diversity. Table S11 reports the estimation results for the engagement at news sharer level. Table S12 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
News | Post | News Post | |
@@
Lifetime | Post Count | |

### v3

@@ Acknowledgements.
| News claim | News post | Claim Post | |
@@
| 0.179 | ✗ | ✗ |
@@
| News lifetime | Post count | |
@@
| 0.481 | 0.188 |
@@
| Anger | ∗∗∗ | ∗∗∗ | ∗∗ |
| Disgust | ∗∗ | ||
| Fear | ∗ | ∗ | |
| Joy | ∗ | ∗∗∗ | |
| Sadness | |||
| Surprise | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✓ | ✓ | ✓ |
| (1) | (2) | (3) | |
| Repost count | Like count | Reply count | |
| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Diversity | ∗∗∗ | ∗∗∗ | ∗∗ |
@@
We provide the cross-correlations to show how the independent variables relate to each other at the news claim and news post levels. Table S5 reports the cross-correlations among the variables in news claims. Table S6 reports the cross-correlations among the variables in news posts.
Additionally, we examine the Variance Inflation Factors (VIFs) among the independent variables to mitigate the concern about multicollinearity. The variance inflation factors are reported in Table S7. We find that the variance inflation factors for the independent variables are all close to one and well below the critical threshold of four, which indicates that multicollinearity is not an issue of our analysis.
@@
Appendix S3 Robustness Check With Original 26 Topics
We adopt the original 26 topics and associated keywords directly from previous research (Chandrasekaran et al., 2020) and repeat the analysis. Table S8 reports the estimation results for the topic diversity. Table S9 reports the estimation results for the engagement at news source sharer level. Table S10 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
| (1) | (2) | (3) | |
| News claim | News post | Claim Post | |
| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗∗∗ | ∗∗∗ | |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗ | ||
| Verified | ∗∗∗ | ||
| AccountAge | |||
| Followers | ∗∗∗ | ||
| Followees | ∗∗∗ | ∗∗∗ | |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✗ | ✓ | ✓ |
| 0.196 | ✗ | ✗ |
| (1) | (2) | |
| News lifetime | Post count | |
| Falsehood | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗ | |
| Diversity | ||
| Falsehood Diversity | ||
| Words | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ |
| Language | ✓ | ✓ |
| MonthYear | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ |
| 0.478 | 0.185 |
| (1) | (2) | (3) | |
| Repost count | Like count | Reply count | |
| Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Falsehood Diversity | |||
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✓ | ✓ | ✓ |
@@
We recalculate topic diversity at the theme level and repeat our analysis. Table S11 reports the estimation results for the topic diversity. Table S12 reports the estimation results for the engagement at news source sharer level. Table S13 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
| News claim | News post | Claim Post | |
@@
| News lifetime | Post count | |
@@
To reduce potential unobserved confounding factors and ensure the robustness of our results, we further conduct propensity score matching for true and false news claims. Specifically, we match true and false news claims based on the variables of emotions and the number of words, with settings including the caliper of 0.3, “common support”, and “noreplacement.” Consequently, we discard 30 false news claims and achieve all biases lower than 0.05. Subsequently, we repeat our analysis based on the matched true and false news claims and their associated posts on X. Table S14 reports the estimation results for the topic diversity. Table S15 reports the estimation results for the engagement at news source sharer level. Table S16 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
| News claim | News post | Claim Post | |
@@
| News lifetime | Post count | |
@@
Appendix S6 Robustness Check With Inclusion of QAnon Conspiracy Theory
We further include QAnon to extend conspiracy theories and repeat our analysis. QAnon is a far-right conspiracy theory alleging that a cabal of politicians and media personalities are engaged in a war to undermine society (Papasavva et al., 2022; Bär et al., 2023; Kim and Kim, 2023). The QAnon discourse has become more diverse and evolved into a global phenomenon during the COVID-19 pandemic (Hoseini et al., 2023). We adopt a list of QAnon-specific keywords from previous research (Sharma et al., 2022a) to identify QAnon conspiracy theory in our dataset. The performance of this keyword list in identifying QAnon discourse has been manually validated, showing high agreement with human assessments (Bär et al., 2023). As a result, we identify 6 QAnon news claims and 302 QAnon news posts. Table S17 reports the estimation results for the topic diversity. Table S18 reports the estimation results for the engagement at news source sharer level. Table S19 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
| (1) | (2) | (3) | |
| News claim | News post | Claim Post | |
| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗∗∗ | ∗∗∗ | |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗∗ | ∗∗∗ | |
| Verified | ∗∗∗ | ||
| AccountAge | |||
| Followers | ∗∗∗ | ||
| Followees | ∗∗∗ | ∗∗∗ | |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✗ | ✓ | ✓ |
| 0.179 | ✗ | ✗ |
| (1) | (2) | |
| News lifetime | Post count | |
| Falsehood | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗∗ | |
| Diversity | ||
| Falsehood Diversity | ||
| Words | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ |
| Language | ✓ | ✓ |
| MonthYear | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ |
| 0.481 | 0.188 |
| (1) | (2) | (3) | |
| Repost count | Like count | Reply count | |
| Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✓ | ✓ | ✓ |
Appendix S7 Robustness Check With Exclusion of Potential Debunking Posts
News posts can also debunk the linked news items. We use keywords of “misinformation,” “misleading,” “fake,” “false,” and “disinformation” to identify potential debunking posts, resulting in news posts. We exclude these potential debunking posts and repeat our analysis. Table S20 reports the estimation results for the topic diversity. Table S21 reports the estimation results for the engagement at news source sharer level. Table S22 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
| (1) | (2) | (3) | |
| News claim | News post | Claim Post | |
| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗∗∗ | ∗∗∗ | |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗ | ∗∗∗ | |
| Verified | ∗∗∗ | ||
| AccountAge | |||
| Followers | ∗∗∗ | ||
| Followees | ∗∗∗ | ∗∗ | |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✗ | ✓ | ✓ |
| 0.173 | ✗ | ✗ |
| (1) | (2) | |
| News lifetime | Post count | |
| Falsehood | ∗∗∗ | ∗∗∗ |
| Conspiracy | ∗ | |
| Diversity | ||
| Falsehood Diversity | ||
| Words | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ |
| Language | ✓ | ✓ |
| MonthYear | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ |
| 0.496 | 0.191 |
| (1) | (2) | (3) | |
| Repost count | Like count | Reply count | |
| Falsehood | ∗∗ | ∗∗∗ | |
| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Falsehood Diversity | ∗∗∗ | ∗ | ∗∗ |
| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗ |
| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
| Emotions | ✓ | ✓ | ✓ |
| Language | ✓ | ✓ | ✓ |
| MonthYear | ✓ | ✓ | ✓ |
| Intercept | ∗∗∗ | ∗∗∗ | |
| News-level REs | ✓ | ✓ | ✓ |
@@
To ensure the generalization of our results across the topic modeling methods, we use BERTopic, a popular unsupervised topic modeling tool, as an alternative method to automatically cluster topics.999https://maartengr.github.io/BERTopic/index.html The BERTopic model generates topic clusters for the whole news claims and posts. We use the topic clusters and their corresponding probabilities to calculate topic diversity again and repeat the analysis. Table S23 reports the estimation results for the topic diversity. Table S24 reports the estimation results for the engagement at news source sharer level. Table S25 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
@@
| News claim | News post | Claim Post | |
@@
| News lifetime | Post count | |
@@
Appendix S9 Additional Analysis on Quote Count
Table S26 reports the full estimation results for quote count.
| Quote count | |
| Falsehood | |
| Conspiracy | ∗∗∗ |
| Diversity | ∗∗∗ |
| Falsehood Diversity | ∗∗∗ |
| Words | ∗∗∗ |
| Media | ∗∗∗ |
| Verified | ∗∗∗ |
| AccountAge | ∗∗∗ |
| Followers | ∗∗∗ |
| Followees | ∗∗∗ |
| Emotions | ✓ |
| Language | ✓ |
| MonthYear | ✓ |
| Intercept | ∗∗∗ |
| News-level REs | ✓ |